OpenStudy (anonymous):

find the equation of a line that passes through (2,8) and is parallel to y=3x+7

Nnesha (nnesha):

parallel lines have same slope

OpenStudy (aaliyahfairgood):

slope intercept form y-y1=m(x-x1) (x,y) =(2,8) m=slope => 3 because it is parallel y-8=3(x-2)

OpenStudy (aaliyahfairgood):

could you solve from there by distributive property

OpenStudy (anonymous):

so would it be y=3x-6

OpenStudy (aaliyahfairgood):

don't forget the 8 y-8=3(x-2)

OpenStudy (aaliyahfairgood):

y-8=3x-6

OpenStudy (anonymous):

y=3x+2

Nnesha (nnesha):

point slope form \[y - y _{1} = m( x -x _{1})\] slope intercept form is y = mx + b

OpenStudy (aaliyahfairgood):

yes! good job :)

OpenStudy (anonymous):

@aaliyahfairgood but I thought the slope had to be 3?

OpenStudy (anonymous):

nvm it is lol, I just had a dumb moment.

OpenStudy (aaliyahfairgood):

y=mx+b y is the line m is the slope m=3 x is multiplied by the slope b is the y intercept

OpenStudy (anonymous):

@aaliyahfairgood would you be able to help me solve this one?

OpenStudy (aaliyahfairgood):

yeah sure

OpenStudy (anonymous):

I think the slope would be -5? but I'm not sure where to go from there.

OpenStudy (aaliyahfairgood):

y-y1=m(x-x1) (x,y) =(7,-1) m =-5/1 (or just-5) because perpendicular lines are negative reciprocals y-7=-5(x-(-1)) negative times negative equal a positive y-7=-5(x+1) y-7=-5x-5

OpenStudy (anonymous):

y=-5x+2?

OpenStudy (aaliyahfairgood):

yes you got it :)

OpenStudy (aaliyahfairgood):

good job